OpenStudy (mic123):

the distance traveled at a constant speed is directly proportional to the time of travle. if olivia traveled 306 miles in 8.5 hours, how many miles will olivia travel in 9.1 hours at the same constant speed? can any one please help me solve this?

OpenStudy (anonymous):

divide 306 by 8.5 to give miles per hour, then multiply by 9.1 hours

OpenStudy (mic123):

thanks so much

OpenStudy (daniellelovee):

@satellite73 can you go to where I tagged you and help me please?

OpenStudy (anonymous):

i.e. \[9.1\times \frac{306}{8.5}\] yw

umerlodhi (umerlodhi):

im gonna say 327.6?

umerlodhi (umerlodhi):

im 95% sure thats the correct answer^^^

OpenStudy (mic123):

could you help with a few more?

umerlodhi (umerlodhi):

sure i'll try

OpenStudy (mic123):

what is the slope of the line that passes through the given points? (3,2) and (5,12)

umerlodhi (umerlodhi):

1 sec

OpenStudy (mic123):

ok

umerlodhi (umerlodhi):

Use the slope formula, the slope of the line through points (x1, y1) and (x2, y2) is (y2 - y1)/(x2 - x1). It doesn't make a difference which point you call (x1, y1) and which you call (x2, y2). In your case, slope = (12 - 2)/(5 - 3) Simplify the fraction

OpenStudy (anonymous):

327.6 looks good to me too http://www.wolframalpha.com/input/?i=9.1*306%2F8.5

OpenStudy (mic123):

isn't it like (x-y) over x1-y1)?

OpenStudy (anonymous):

no it isn't from 3 to 5 is two right from 2 to 12 is ten up right 2, up 10, slope is \(\frac{10}{2}\)

OpenStudy (mic123):

oh ok

OpenStudy (mic123):

so it'll be 5?